OpenStudy (anonymous):

finding simplest form > A meat company cuts slices of pepperoni with a thickness of 3/8 inch. If the stick of pepperoni is 12 inches long, how many slices can be made from one stick?

OpenStudy (anonymous):

\[12 = \frac{ 96 }{ 8 } So \frac{ 96 }{ 8}*\frac{ 8 }{ 3 } = 32\]

OpenStudy (anonymous):

yay I was right ill give u a medal